# MinD-3D: Reconstruct High-quality 3D objects in Human Brain

### v1

@@ 5.2 Neuro-Fusion Encoder
Drawing from the potent Masked Image Model (MIM) [16, 26], we adopt an auto-encoder structure to extract meaningful fMRI features, trained on the UKB dataset [23]. Our encoder, termed the Neuro-Fusion Encoder (NFE), utilizes the encoder of LEA [26] as $E_{f}$ with a Feature Aggregation module $\mathcal{FA}$ to transform fMRI frames into an embedding compatible with the CLIP [28] ViT-B/32 model’s space.
@@
where n is the number of selected video frames. Then we utilize the CLIP loss to ensure the alignment between the fMRI features $\mathbf{c}_{f}$ and the visual features $\mathbf{c}_{v}$:
@@
During training, we optimize the whole Neuro-Fusion Encoder with $E_{LEA}$ loading the pre-trained weights and the CLIP vision encoder frozen. It’s worth noting that the CLIP vision encoder $E_{v}$ and images $\{\mathbf{V}\}$ are just used for training, we will drop it during inference. By adopting this approach, we will get the fMRI features $\mathbf{c}_{f}$ and embedding $\mathbf{F}_{emb}$ aligned with the visual space. The fMRI features $\mathbf{c}_{f}$ will be used as the conditional information for the feature bridge diffusion model and fMRI embedding $\mathbf{F}_{emb}$ will be used in the latent adapted decoder.
@@
LEA-3D[26] 0.787 0.371 0.353 0.527 0.562 0.575 0.624 0.372
fMRI-PTE-3D[27] 0.815 0.392 0.375 0.433 0.694 0.327 0.519 0.135
w/o Both 0.784 0.364 0.340 0.536 0.618 0.498 0.649 0.246
w/o Diffusion 0.802 0.382 0.372 0.509 0.641 0.367 0.546 0.175
w/o Contrastive 0.822 0.409 0.398 0.453 0.674 0.236 0.504 0.095
MinD-3D (full) 0.835 0.425 0.411 0.419 0.704 0.214 0.480 0.065
Table 1: Performance Comparison on the fMRI-Shape. We report average metrics for all subjects. LEA-3D and fMRI-PTE-3D represent variants of LEA and fMRI-PTE, respectively. The following 3 baselines are degenerate versions of MinD-3D.
5.3 Feature Bridge Diffusion Model
This phrase details the Neural-Visual Synthesis process using a diffusion model, named here as the Feature Bridging Diffusion Model (FBDM), following the design outlined in [29]. This module’s primary function is to facilitate the generation of the target visual latent feature $\hat{\mathbf{c}_{v}}$ from the corresponding fMRI latent feature $\mathbf{c}_{f}$.
In the forward diffusion process, we treat the visual latent feature $\mathbf{c}_{v}$ as $\mathbf{x}_{0}$ and introduce Gaussian noise into it across 100 timesteps. This approach modifies the standard 1000-timestep model to align with the settings in Mind-eye [31]:
where, $\mathbf{\epsilon}$ is the Gaussian noise. During the reverse diffusion process, the module will learn to predict and then reverse the noise introduced during the forward process. The goal is to accurately estimate the noise $\hat{\mathbf{\epsilon}}_{t}$ at each timestep, conditioned on the fMRI latent feature $\mathbf{c}_{f}$:
$\hat{\mathbf{\epsilon}}_{t}=FBDM(\mathbf{x}_{t},t,\mathbf{c}_{f})$
The fMRI latent feature $\mathbf{c}_{f}$ is mapped to the intermediate layers of the FBDM through a cross-attention operation:
CrossAttention$(Q,K,V)=\operatorname{softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)\cdot V$, with
$Q=W_{Q}^{(i)}\varphi_{i}\left(x_{t}\right),K=W_{K}^{(i)}\tau_{\theta}(\mathbf{c}_{f}),V=W_{V}^{(i)}\tau_{\theta}(\mathbf{c}_{f}).$
Here, $\tau_{\theta}$ is the fMRI latent feature with a suitable dimension projector, $\varphi_{i}\left(x_{t}\right)$ denotes intermediate values of the Decoder and $W_{Q}^{(i)}$, $W_{K}^{(i)},W_{V}^{(i)}$ are projector matrices with learnable parameters.
So the training objective of FBDM is to minimize the difference between the actual and predicted noise:
Post-training, the visual latent feature $\mathbf{c}_{v}$ will be dropped and FBDM initiates a generative phase, starting with pure noise and progressively applying the reverse diffusion process to produce the predicted visual latent feature $\hat{\mathbf{c}_{v}}$. This phase crucially maps the aligned fMRI latent features $\mathbf{c}_{f}$ to the visual space and generates the corresponding visual latent features $\hat{\mathbf{c}_{v}}$, which is used as the basic conditional information for the latent adapted decoder.
5.4 Latent Adapted Decoder
Latent Adapted Decoder is based on Argus [39], a cutting-edge 3D shape generator. It will generate the VQ latent representation of the target 3D object based on the conditional information we provide via a generative auto-regressive transformer and then get the 3D objects through the VQ decoder. Here, we design an adapter layer in the transformer block to modify the module using the fMRI embeddings $\mathbf{F}_{emb}$.
Concretely, to directly incorporate the fMRI feature into the 3D generation process, we adapt the Multihead Self-Attention (MSA) in the standard transformer block into a double attention layer. This layer consists of the original MSA and an adapter layer. The adapter layer calculates multi-head cross-attention using the intermediate values of the module. We then obtain the adapted result using a learnable parameter $\lambda$ as the coefficient:
$x=\lambda\cdot\textsc{CA}(Q,K_{f},V_{f})+(1-\lambda)\cdot\textsc{SA}(Q,K,V)$
where CA and SA indicate cross-attention and self-attention separately; $Q,K,V$ denotes the representation of 3d latent features with a Linear projection; $K_{f},V_{f}$ are the representation of fMRI embeddings with a linear projection. The adapter layer is added every T transformer block.
Our decoder computes the VQ latent vector for the target object:
$p(\mathbf{z})=\prod_{i=1}^{m}p\left(\mathbf{z}_{i}\mid\hat{\mathbf{c}_{v}},\mathbf{F}_{emb},\mathbf{z}_{<i}\right)$
Finally, the 3D VQ-decoder decodes this latent vector $\mathbf{z}$ to reconstruct the final point clouds $\mathbf{P}$:
$\{\mathbf{P}\}=D(\mathbf{z})$
The decoder’s optimization objective is defined by the negative log-likelihood loss:
During training, we train the module separately using pre-training weights, after training the neuro-fusion encoder and the feature bridge diffusion model.
6 Expriment
@@ 6.1 Metrics
In order to evaluate the performance of the new task and our model comprehensively, we adopt the following metrics in two aspects, semantic-level and structure-level:
Semantic Level. To evaluate the semantic quality of our model, we adopt standard metrics used in previous 2D fMRI literature [24, 4, 5, 25, 22], namely pixel-level correlation (PixCorr) and N-way top-K accuracy. These two metrics are calculated by comparing the reconstructed images with the ground truth (GT) images. In this process, we render both the reconstructed objects and the GT objects into images separately, calculate the metrics for each frame, and then derive the average value as the final metric numbers.
Structure Level. Besides evaluating the semantic quality of our model, it is also essential to assess its performance in terms of spatial structure, which we do in both 3D and 2D. In 3D, we utilize common metrics used in 3D reconstruction: Chamfer Distance (CD), Hausdorff Distance (HD), and Earth Mover’s Distance (EMD). For the 2D evaluation, we render the 3D objects onto a 2D plane and compute the frame-averaged pixel-level correlation, Learned Perceptual Image Patch Similarity (LPIPS), and Structural Similarity (SSIM).
6.2 Implementation Details.
As described in Section 3, each data pair in the fMRI-Shape dataset consists of 10 fMRI frames and the corresponding 192 images. To maximize the utilization of this dataset, we randomly select 6 fMRI frames for training, which also serves as a form of data augmentation. These 6 frames, initially 1023 × 2514 2D fMRI images, are resized to 256 × 256. For the image data, we set the number of selected video frames (n) to 4, meaning we randomly choose 4 frames for subsequent training, and each of these images is resized to 224 × 224. During inference, we select only the middle 6 fMRI frames and do not input any additional images. In all experiments, we train the model using data from all 3 subjects and then calculate the metrics for each subject.
Regarding the model structure, our neuro-fusion encoder comprises 24 transformer layers with 1024 feature dimensions and 16 self-attention heads. Utilizing the CLIP ViT-B/32 model, our Feature Aggregation module maps features to a 512-dimensional space. Consequently, the feature bridge diffusion model is configured with a feature dimension of 512 for cross-attention and 8 self-attention heads, with the number of diffusion channels set to 224. The latent adapted decoder is designed with 32 transformer layers, 3072 feature dimensions, and 16 self-attention heads. We set T=8, which means that we add an adapter layer every 8 transformer blocks.
In terms of model training, the first stage involves optimizing both the neuro-fusion encoder and the feature bridge diffusion model. We initialize the neuro-fusion encoder with pre-trained weights, and the feature bridge diffusion model is trained from scratch. In the second stage, we fix the parameters of the first two stages and focus on optimizing the latent adapted decoder. We use the AdamW optimizer with a learning rate of 5e-5 for all experiments and train the model with a batch size of 128 for all subjects.
@@ 6.3 Experimental Results
As the first work to model 3D imaging in the human brain, we compare our comprehensive model, MinD-3D, with several baselines, including adaptations of LEA and fMRI-PTE using the same 3D decoder as in MinD-3D. Tab. 1 presents the averaged metrics at both the Structural and Semantic Levels for all subjects. MinD-3D demonstrates superior performance across all metrics: achieving approximately 83.5% accuracy in 2-way-top-1, 42.5% accuracy in 10-way-top-1, a pixel-level correlation of 0.411, and 0.419 in LPIPS. These results indicate that our model not only generates objects with high semantic accuracy but also excels in structural similarity, outperforming baseline methods.
@@
6.4 Analytical and Ablation Study
We further conduct some analytical and ablation studies to evaluate the effectiveness of our model.
Figure 5: Quality of the extracted features. We map the features to brain regions via ridge regression modeling, with voxel-wise Pearson’s correlation coefficients computed between them. The visual function ROI is highlighted in red, indicating correlations in the range of 0 to 0.8. In contrast, other regions are depicted in blue, representing correlations within a range of 0 to 0.5.
Analysis of feature extracted. To understand our feature extractor from a biological perspective, following [14, 37, 35, 24], we employ a linear encoding model here to project the extracted fMRI feature $c_{f}$ to brain activity. For this purpose, we select subject 1 as our representative. Specifically, we employ ridge regression to learn the mapping relationship between $c_{f}$ and fMRI signals in the training set. We then map $c_{f}$ extracted from the test set to the fMRI space and calculate the Pearson correlation coefficient. As illustrated in Fig. 5, we use red to indicate the positive correlation between visual ROIs. For other regional positive correlations, we use blue. A high positive correlation is evident between $c_{f}$ and visual ROIs, indicating that we have successfully extracted valid features. Additionally, we conducted a similar analysis with the visual latent feature $c_{v}$ and visualized its positive correlations. Notably, the correlation of $c_{f}$ is superior to that of $c_{v}$. Furthermore, we calculated the correlation between $c_{f}$ and $c_{v}$, which is very high. This high correlation suggests that our comparative learning is effective and that the fMRI feature space and image space are well-aligned.
Ablation study of fMRI frames. During training, we randomly select 6 fMRI frames, denoted as $F_{n}$, as the input to our model. To explore the impact of different $F_{n}$ values on our model’s performance, we conducted an assessment across various $F_{n}$ values. We report and compare average results for metrics such as 2-way-top-1, 10-way-top-1, SSIM, and CD, focusing on the performance across three subjects. As shown in Tab. 2, $F_{n}=6$ emerges as the optimal choice. This operation not only acts as data augmentation but also, due to the inherent noise in fMRI signals, choosing the middle portion for inference helps in reducing the introduction of noise.
@@ 7 Conclusion
$F_{n}$ Semantic-level Structure-Leve
@@
1 0.674 0.291 0.472 0.853
3 0.762 0.365 0.604 0.526
6 0.835 0.425 0.704 0.214
8 0.829 0.401 0.702 0.225
10 0.818 0.379 0.684 0.273
Table 2: Ablation results for the fMRI frames. It indicates that $F_{n}=6$ is the best choice for MinD-3D.
In this paper, for the first time, we introduce the task of 3D reconstruction based on fMRI signals, accompanied by a dedicated dataset. Technically, we propose a novel three-stage framework tailored for this groundbreaking task. In the initial stage, our model adeptly extracts features from fMRI frames. Subsequently, in the second stage, it employs a diffusion module to translate these fMRI features into the image domain. The final stage involves the transformation of these image features into 3D models using an advanced 3D generation model. Our experimental results and analyses demonstrate the efficacy of our model in accurately extracting fMRI features and converting them into their corresponding 3D objects. This work not only establishes a new domain in neuroimaging and 3D reconstruction but also opens avenues for further research in understanding and visualizing neural representations in 3D vision.
Appendix A More Implementation Details

### v3

@@ 5 Method
Figure 4: Overview of the MinD-3D Framework. Our approach combines a Neuro-Fusion Encoder for extracting features from fMRI frames, a Feature Bridge Diffusion Model for generating visual features from these fMRI signals, and a Latent Adapted Decoder based on the Argus 3D shape generator for reconstructing 3D objects. This integrated system effectively aligns and translates brain signals into accurate 3D visual representations. Note that the CLIP encoder is only for training the model, while not used for inference.
@@ 5.2 Neuro-Fusion Encoder
Drawing from the potent Masked Image Model (MIM) [17, 13, 34], we adopt an auto-encoder structure to extract meaningful fMRI features, trained on the UKB dataset [30]. Our encoder, termed the Neuro-Fusion Encoder (NFE), employs the encoder of LEA [34] as $E_{f}$, along with a Feature Aggregation module ($\mathcal{FA}$), to transform fMRI frames into embeddings. Inspired by contrastive learning [37, 9, 57], we then align these embeddings with the vision space through CLIP [37].
@@
where n is the number of selected video frames. Then we utilize the CLIP loss to ensure the alignment between fMRI features $\mathbf{c}_{f}$ and visual features $\mathbf{c}_{v}$:
@@
During training, we optimize the neuro-fusion encoder initialized with pre-trained weights, while the CLIP vision encoder remains frozen. It’s worth noting that the CLIP vision encoder $E_{v}$ and images $\{\mathbf{V}\}$ are just used for training, we will drop it during inference. By adopting this approach, we will get the fMRI features $\mathbf{c}_{f}$ and embedding $\mathbf{F}_{emb}$ aligned with the visual space. The fMRI features $\mathbf{c}_{f}$ will be used as the conditional information for the feature bridge diffusion model.
5.3 Feature Bridge Diffusion Model
This section describes the Neural-Visual Synthesis process via a transformer-based diffusion model, herein referred to as the Feature Bridge Diffusion Model (FBDM), inspired by the hierarchical design proposed in [38]. The FBDM’s primary role is to bridge fMRI latent features $\mathbf{c}_{f}$ with their visual counterparts $\hat{\mathbf{c}_{v}}$, effectively transforming neurological signals into visual representations. In the forward diffusion process, we treat the visual latent feature $\mathbf{c}_{v}$ as $\mathbf{x}_{0}$ and introduce Gaussian noise into it across 100 timesteps. This approach modifies the standard 1000-timestep model to align with the settings in Mind-eye [41]:
During the reverse diffusion process, FBDM is tasked with predicting and reversing the introduced noise to reconstruct the latent visual feature. This is achieved by estimating the noise $\hat{\mathbf{\epsilon}}_{t}$ at each timestep, conditioned on the fMRI latent feature $\mathbf{c}_{f}$:
$\hat{\mathbf{\epsilon}}_{t}=FBDM(\mathbf{x}_{t},t,\mathbf{c}_{f})$
Differing from U-Net based models, our approach utilizes a transformer to leverage the self-attention mechanism for integrating the conditional information $\mathbf{c}_{f}$. This is done by embedding the timestep $t$, the noise-perturbed feature $\mathbf{x}_{t}$, and the conditional fMRI feature $\mathbf{c}_{f}$ into a unified representation, further augmented with a learnable token to capture the target vision features accurately. The objective of FBDM training is to minimize the discrepancy between the actual noise and its prediction:
Upon training completion, the model discards the initial visual latent feature $\mathbf{c}_{v}$ and enters a generative phase. It begins with pure noise and progressively refines this through the reverse diffusion process to generate the predicted visual latent feature $\hat{\mathbf{c}_{v}}$. This process effectively translates the aligned fMRI latent features $\mathbf{c}_{f}$ into visual representations, serving as essential conditional information for the subsequent latent-adapted decoding phase.
@@ 5.3 Feature Bridge Diffusion Model
5.4 Latent Adapted Decoder
The Latent Adapted Decoder (LAD) leverages Argus [35], a state-of-the-art 3D mesh generation method, as its backbone. This component is tasked with generating the Vector Quantized (VQ) latent representation of target 3D objects from conditional information processed through a generative auto-regressive transformer. The resulting 3D objects are then reconstructed using the VQ decoder.
To tailor Argus for processing fMRI input and to retain its generative efficiency, we employ adapter tuning during training. This involves integrating the adapter layers within the transformer block. Following classic methods [20, 53], we apply two feed-forward linear layers to effectively adapt the self-attention outputs. The adapter layer is systematically inserted at every $T$ transformer blocks, ensuring that the model remains adaptable and efficient. The decoder formulates the VQ latent vector for the target object as follows:
$p(\mathbf{z})=\prod_{i=1}^{m}p\left(\mathbf{z}_{i}\mid\hat{\mathbf{c}_{v}},\mathbf{z}_{<i}\right)$
Subsequently, the 3D VQ-decoder decodes this latent vector $\mathbf{z}$ into the final 3D mesh representation $\Psi$:
$\Psi=D(\mathbf{z})$
The decoder aims to minimize the negative log-likelihood loss during optimization:
For training, the decoder is fine-tuned from pre-trained weights, with the trained neuro-fusion encoder and the feature bridge diffusion model. This strategy preserves Argus’s generative capabilities while adapting to new inputs and minimizing the computational overhead associated with training.
Table 2: Performance Comparison on fMRI-Shape. We report the average metrics for each subject trained and tested on their own data in the core part. LEA-3D and fMRI-PTE-3D represent variants of LEA and fMRI-PTE, respectively. The following three baselines are ablation studies for MinD-3D.
@@ 5.4 Latent Adapted Decoder
LEA-3D[34] 0.787 0.371 0.527 0.562 4.229 2.291 5.347
fMRI-PTE-3D[36] 0.815 0.392 0.433 0.694 3.571 1.992 4.621
w/o Both 0.789 0.367 0.479 0.616 3.694 2.205 5.073
w/o Diffusion 0.801 0.385 0.423 0.669 3.526 2.071 4.625
w/o Contrastive 0.823 0.419 0.319 0.701 3.315 1.826 4.027
MinD-3D (full) 0.839 0.432 0.230 0.734 3.157 1.742 3.833
6 Experiments
@@ 6.1 Metrics
To more effectively evaluate the performance of our models on the new task, we utilize metrics in two aspects: semantic-level and structure-level.
Semantic Level. To evaluate the semantic quality of our model, we adopt standard metrics used in previous 2D fMRI literature [6, 7, 31, 29, 45], namely, N-way top-K accuracy. We calculate 2-way-top-1 and 10-way-top-1 accuracy as shown in Tab. 2. Additionally, we calculate the Learned Perceptual Image Patch Similarity (LPIPS) [55] to assess the perceptual quality of semantics. In practice, these two metrics are calculated by comparing the reconstructed images with the ground truth (GT) images. In this process, we render both the reconstructed and GT objects into images separately at every 60 degrees, calculate the metrics for each frame, and then derive the average value as the final metric numbers.
Structure Level. In addition to the semantic quality evaluation, it is crucial to assess the performance of our model in capturing geometrical structures. We utilize common metrics employed in 3D reconstruction [28, 52, 35]: Fréchet Point Cloud Distance (FPD, adjusted by a factor of ×$\times 10^{-1}$), Chamfer Distance (CD, scaled by ×$\times 10^{2}$), and Earth Mover’s Distance (EMD, also scaled by ×$\times 10^{2}$). We sample point clouds from the GT and reconstructed meshes to calculate these three metrics. Furthermore, we compute the frame-averaged Structural Similarity Index (SSIM) using the same rendering method to assess similarity from each viewpoint.
6.2 Implementation Details
As detailed in Sec. 3, each data pair in the fMRI-Shape dataset is composed of 10 fMRI frames and 192 corresponding images. To maximize dataset utilization and introduce data augmentation, we randomly select 6 fMRI frames for training. These frames, originally 1023 × 2514 2D fMRI images, are resized to 256 × 256. For the image data, 8 frames are randomly selected for training, with each image resized to 224 × 224. During inference, only the middle 6 fMRI frames are used, without incorporating additional images. Regarding the architecture, we set $T=4$ to incorporate adapter layers.
The model training proceeds in two stages. Initially, we focus on optimizing the neuro-fusion encoder and the feature bridge diffusion model. The neuro-fusion encoder is initialized with pre-trained weights, while the feature bridge diffusion model undergoes training from scratch, requiring approximately 2.5 days on a single A100 GPU. In the second stage, we optimize all parameters, including those of the adapters, and freeze the parameters in Argus. This stage is completed in about one day on a single A100 GPU. Details on the hyperparameters of the model are provided in the supplementary material.
@@ 6.3 Experimental Results
As the pioneering effort in modeling 3D imaging within the human brain, which involves various regions of the brain, direct comparisons with existing models are not feasible. To mitigate the training costs, we have adapted LEA and fMRI-PTE models, which were trained on the same vision ROIs, utilizing the identical 3D decoder as employed in MinD-3D. These adaptations serve as the baselines in this study, facilitating a more contextual and fair comparison within the novel domain we are exploring. Tab. 2 presents the averaged metrics at both the Structural and Semantic Levels for all subjects. MinD-3D demonstrates superior performance across all metrics: achieving approximately 83.9% accuracy in 2-way-top-1, 43.2% accuracy in 10-way-top-1, and scores of 0.734, 0.230, 3.157, 1.742, and 3.833 in other metrics. These results indicate that our model not only generates objects with high semantic accuracy but also excels in structural similarity, outperforming baseline methods. Tab. 2 also proves the effectiveness of the diffusion model and the contrastive learning module, both of which play important roles in our MinD-3D.
@@
6.4 Further Analysis
To effectively utilize a subset of the fMRI-Shape dataset and assess the generalization capabilities of our proposed MinD-3D model, we conducted two Out-Of-Distribution (OOD) experiments under challenging settings: 1) Across-Person Testing (APT): In APT, we evaluate our model, which was trained only on Subject 1, using the data from Subject 9. We compare the results with the baselines and report the metrics in Tab. 3. 2) Across-Person & Across-Class Testing (APACT): In APACT, we similarly evaluate our model, trained solely on Subject 1, with the data from Subject 11. We also compare with the baselines and report the metrics in Tab. 3.
Although the performance in these OOD scenarios does not match the In-Distribution (ID) results, which is to be expected for such an extremely challenging task and given the substantial individual differences and domain gaps, our method still outperforms the baselines. This achievement not only underscores the robustness of MinD-3D but also establishes a new benchmark for the community to build upon.
@@ 7 Conclusion
Table 3: Quantitative results of APT and APACT. We use the model trained on Subject 1 and compare metrics for Subjects 9 and 11 separately.
Methods APT APACT
@@
LEA-3D 5.362 3.627 6.174 6.958 4.944 8.107
fMRI-PTE-3D 4.501 2.956 5.772 6.261 4.570 7.843
MinD-3D 3.838 2.415 5.117 5.689 4.181 7.194
In this paper, we introduce the innovative task of Recon3DMind and its accompanying large-scale dataset, fMRI-Shape, across various settings for the first time. Technologically, we develop a novel three-stage framework that involves more brain regions, including those associated with 3D vision in humans, specifically designed for this task. This approach sets a new benchmark in the field and proves the feasibility of the task. Initially, our model proficiently extracts features from fMRI frames. It then utilizes a diffusion module in the second stage to transition these fMRI features into the vision domain. The culmination of this process is the transformation of vision features into 3D models by an advanced 3D generation model in the final stage. Our comprehensive experimental results and analyses affirm the model’s effectiveness in precisely extracting fMRI features and converting them into their corresponding 3D objects. This pioneering work not only carves out a new niche in neuroimaging and 3D reconstruction but also paves the way for future research aimed at a deeper understanding and visualization of neural representations in 3D vision.
Acknowledgements
The computations in this research were performed using the CFFF platform of Fudan University.
Appendix 0.A FMRI-Shape Details
0.A.1 Visualization of fMRI Across Different Subjects
To more effectively demonstrate brain activation patterns and showcase our fMRI-Shape dataset, we analyze and visualize the responses to five distinct objects across six different subjects in Fig. 0.A.1. (Note: Only voxels with activation levels exceeding the 50th percentile are displayed.)
Figure 0.A.1: Individual differences in brain activation patterns within the fMRI-Shape.
Fig. 0.A.1 also reveals significant individual differences in brain activation across subjects, more so than the variations in responses to different objects. This observation underscores the challenges inherent in the settings of AP and APAC.
0.A.2 Visualization of the Core Component
To illustrate the diversity of fMRI-Shape, Fig. 0.A.2 shows the core part, organized with five objects of training data and two objects of testing data each category.
0.A.3 Visualization of the APAC Set
The APAC Set, which includes additional objects not existing in the Core parts, is showcased in Fig. 0.A.3, with one object displayed for each category.
Figure 0.A.2: Samples of Core component in fMRI-Shape. We show five objects of training data and two objects of testing data per category.
Figure 0.A.3: Samples of APAC Set in fMRI-Shape. One object is shown per category.
Appendix 0.B Analysis of Visualizations
0.B.1 Visualizations from Different Subjects
We display several samples from 5 subjects in Fig. 0.B.1 to demonstrate the effectiveness of our MinD-3D model. Due to the varying responses of different subjects to different objects, there are some variations in the results. Overall, our model accurately reconstructs both the category and characteristics of the objects. The experiment validates the feasibility of our proposed task, Recon3DMind, and the accompanying dataset, fMRI-Shape, designed for this task.
0.B.2 Visualizations of AP & APAC Testing
We also showcase the reconstructed objects from AP & APAC testing in Fig. 0.B.2. As highlighted in Fig. 0.A.1, individual differences significantly affect the outcomes. Our AP & APAC Testing empirically confirms this. Despite the task’s high difficulty, MinD-3D successfully recovers the basic shapes of the objects, establishing a baseline method for the community.
Figure 0.B.1: Samples from different subjects.
Figure 0.B.2: Visualization of AP & APAC testing. AP Testing trains on Subject 1 and tests on Subject 9. APAC Testing trains on Subject 1 and tests on Subject 11.
0.B.3 Fail Cases
We present some failure cases in Fig. 0.B.3. Although the reconstruction does not always accurately capture the entire object, it still manages to retain the shape or characteristics of the objects.
Figure 0.B.3: Fail cases.
Appendix 0.C More Implementation Details